OpenStudy (anonymous):

The value of a two digit number is twice as large as the sum of its digits. If the digits were reversed, the resulting number would be 9 less than 5 times he original number. What is the original number?

OpenStudy (anonymous):

10.a+b=N a+b=N/2 10.b+a=5N-9 solve it

OpenStudy (ikram002p):

a+10b= 2 (a+b) 10a+b=5(10b+a)-9